Determine the value of a and b so that lim(x→0) (asin2x-bsinx/x³)=1


1
Expert's answer
2022-01-17T18:22:47-0500
"\lim\limits_{x\to 0}\dfrac{a\sin(2x)-b\sin x}{x^3}=[\dfrac{0}{0}]"

"=\lim\limits_{x\to 0}\dfrac{(a\sin(2x)-b\sin x)'}{(x^3)'}"

"=\lim\limits_{x\to 0}\dfrac{2a\cos(2x)-b\cos x}{3x^2}=[\dfrac{0}{0}]"


"2a\cos(2(0))-b\cos (0)=0=>2a=b"

"\lim\limits_{x\to 0}\dfrac{b\cos(2x)-b\cos x}{3x^2}=[\dfrac{0}{0}]"

"=\lim\limits_{x\to 0}\dfrac{(b\cos(2x)-b\cos x)'}{(3x^2)'}"

"=\lim\limits_{x\to 0}\dfrac{-2b\sin(2x)+b\sin x}{6x}=[\dfrac{0}{0}]"


"=\lim\limits_{x\to 0}\dfrac{(-2b\sin(2x)+b\sin x)'}{(6x)'}"

"=\lim\limits_{x\to 0}\dfrac{-4b\cos(2x)+b\cos x}{6}"

"=\dfrac{-4b\cos(2(0))+b\cos (0)}{6}"

"=-\dfrac{b}{2}=1=>b=-2, a=-1"

"\lim\limits_{x\to 0}\dfrac{-\sin(2x)+2\sin x}{x^3}=1"

"a=-1, b=-2"